# Impact of Synthetic Images on Morphing Attack Detection Using a Siamese Network

### 14 Mar 2024 v1

@@ 3.2 SOTA Databases
Figure 1: Example of images from all the databases. a) SDD: Left to Right: Bona fide (Subject-1) synthetic, Morph Synthetics (Subject-2), MIPGAN-I, MIPGAN-II, OpenCV-Morpher, WebMorpher. B) FERET: Left to right: Bona fide (Subject-1), Bona fide (Subject-2), FaceMorpher, FaceFusion, OpenCV-Morpher, UBO-Morpher. c) FRGCv2: Left to right: Bona fide (Subject-1), Bona fide (Subject-2), FaceMorpher, FaceFusion, OpenCV-Morpher, UBO-Morpher. d) FRLL: Left to Right: Bona fide (Subject-1), Bona fide (Subject-2), AMSL, FaceMorpher, OpenCV-Morpher, and StyleGAN.
@@ 5.1 Exp1 - Train SYN-MAD/TEST SDD-Bechmark
For the tests a), b) and c), 23,000 bona fide and 13,000 morphed synthetic images were used to train. For the validation set, 4,000 (2,000 bona fide and 2,000 synthetics) images were used.
@@
OpenCV 27.54 50.49 38.24 22.06 33.59 49.80 (90037) Trained:SDD Train-Set Test: Benchmark Test-Set Syntheticss and Landmark Compared with E-CBAM@VCMI
@@
OpenCV N/A 16.66 28.92 54.51 (134122) Trained: Train-SOTA No synthetrics images Test: Benchmark Test-set
@@
OpenCV 5.69 1.96 1.47 6.37 1.90 0.84 Trained: Mix (synthetics plus SOTA) Test: Benchmark Test-set Compared with MorphHRNet
@@ 5.2 Exp2 - Train SYN-MAD/TEST SOTA
The best results from experiment 1 were used for training the EfficientNetB0 again, but now only using the official training set of the SYN-MAD database release for the testing process was evaluated with SOTA: AMSL, FERET, FRGC and FRLL databases. 23,000 bona fide and 13,000 morphed synthetics images were used to train. For the validation set, 4,000 (2,000 bona fide and 2,000 synthetics) images were used. The best $lr$ was $41e-$ 4. Table 2 shows a summary with the best results for experiment 2. This experiment shows that synthetic images can not reach generalisation capabilities, and the performance is worse when compared with digital images based on landmark morphing methods. The EER obtained was 29.12%. Table 2 shows the summary results.
@@ 5.3 Exp3 - Train SOTA/TEST SDD-Benchmark
For experiment 3, the best results from experiment 1 (EfficientNetB0) also was used for training again, but now used the FERET, FRGC and FRLL images for training and tested in the official test set of the SDD-Benchmark database. The $lr$ was set up to $55e-$ 5. This experiment can infer with a low error the synthetics images even when no synthetics images were used in the training set. The EER, BPCER10 and BPCER20 for SDD-test reached 3.49%, 1.96% and 2.94%, respectively. The results for a cross-set with AMSL reached 2.02% and 3.29%. Table 2 shows a summary with the best results for experiment 3.
@@ 5.4 Exp4 - Train Mix/TEST SDD-Benchmark
For Experiment 4, the synthetic images from the SYN-MAD, FERET, and FRGCv2 were mixed by increasing the number of images available for training. The SDD-test benchmark was used as a test set. This experiment allows us to compare our proposed method with the best results of the IJCB 2022 competition obtained by the MorphHRNet team (Table 2). The best results were obtained when at least 20% of the images belonged to digital faces (no synthetics). There were 23.000 synthetic face images plus 5,000 FERET and FRGC digital bona fide images and 13,000 morph synthetics images. For the validation set, 2,000 bona fide and 2,000 morphs were used. Table 2 shows the summary results. Figure 5 shows an individual analysis of the train-mix database evaluated with SDD benchmark datasets break for each morphing tool separately. Morphed images created by WebMorpher tools have been identified as the most difficult to detect in comparison with FaceMorpher, MIPGAN-I, MIPGAN-II, and OpenCV-Morpher. Figure 5 show the improvements reached when the 20% SOTA database is included in the pure synthetics database. The BPCER10 and BPCER20 for SDD-test reached 0.98% and 5.39%, respectively.

### 20 Mar 2024 v2

@@ 3.2 SOTA Databases
Figure 1: Example of images from all the databases. a) SDD: Left to Right: Bona fide (Subject-1) synthetic, Morph Synthetic (Subject-2), MIPGAN-I, MIPGAN-II, OpenCV-Morpher, WebMorpher. B) FERET: Left to right: Bona fide (Subject-1), Bona fide (Subject-2), FaceMorpher, FaceFusion, OpenCV-Morpher, UBO-Morpher. c) FRGCv2: Left to right: Bona fide (Subject-1), Bona fide (Subject-2), FaceMorpher, FaceFusion, OpenCV-Morpher, UBO-Morpher. d) FRLL: Left to Right: Bona fide (Subject-1), Bona fide (Subject-2), AMSL, FaceMorpher, OpenCV-Morpher, and StyleGAN.
@@ 5.1 Exp1 - Train SYN-MAD/TEST SDD-Bechmark
For the tests a), b) and c), 23,000 bona fide and 13,000 morphed synthetic images were used to train. For the validation set, 4,000 (2,000 bona fide and 2,000 synthetic) images were used.
@@
OpenCV 27.54 50.49 38.24 22.06 33.59 49.80 (90037) Trained:SDD Train-Set Test: Benchmark Test-Set Synthetic and Landmark Compared with E-CBAM@VCMI
@@
OpenCV N/A 16.66 28.92 54.51 (134122) Trained: Train-SOTA No synthetic images Test: Benchmark Test-set
@@
OpenCV 5.69 1.96 1.47 6.37 1.90 0.84 Trained: Mix (synthetic plus SOTA) Test: Benchmark Test-set Compared with MorphHRNet
@@ 5.2 Exp2 - Train SYN-MAD/TEST SOTA
The best results from experiment 1 were used for training the EfficientNetB0 again, but now only using the official training set of the SYN-MAD database release for the testing process was evaluated with SOTA: AMSL, FERET, FRGC and FRLL databases. 23,000 bona fide and 13,000 morphed synthetic images were used to train. For the validation set, 4,000 (2,000 bona fide and 2,000 synthetics) images were used. The best $lr$ was $41e-$ 4. Table 2 shows a summary with the best results for experiment 2. This experiment shows that synthetic images can not reach generalisation capabilities, and the performance is worse when compared with digital images based on landmark morphing methods. The EER obtained was 29.12%. Table 2 shows the summary results.
@@ 5.3 Exp3 - Train SOTA/TEST SDD-Benchmark
For experiment 3, the best results from experiment 1 (EfficientNetB0) also was used for training again, but now used the FERET, FRGC and FRLL images for training and tested in the official test set of the SDD-Benchmark database. The $lr$ was set up to $55e-$ 5. This experiment can infer with a low error the synthetic images even when no synthetic images were used in the training set. The EER, BPCER10 and BPCER20 for SDD-test reached 3.49%, 1.96% and 2.94%, respectively. The results for a cross-set with AMSL reached 2.02% and 3.29%. Table 2 shows a summary with the best results for experiment 3.
@@ 5.4 Exp4 - Train Mix/TEST SDD-Benchmark
For Experiment 4, the synthetic images from the SYN-MAD, FERET, and FRGCv2 were mixed by increasing the number of images available for training. The SDD-test benchmark was used as a test set. This experiment allows us to compare our proposed method with the best results of the IJCB 2022 competition obtained by the MorphHRNet team (Table 2). The best results were obtained when at least 20% of the images belonged to digital faces (no synthetic). There were 23.000 synthetic face images plus 5,000 FERET and FRGC digital bona fide images and 13,000 morph synthetic images. For the validation set, 2,000 bona fide and 2,000 morphs were used. Table 2 shows the summary results. Figure 5 shows an individual analysis of the train-mix database evaluated with SDD benchmark datasets break for each morphing tool separately. Morphed images created by WebMorpher tools have been identified as the most difficult to detect in comparison with FaceMorpher, MIPGAN-I, MIPGAN-II, and OpenCV-Morpher. Figure 5 show the improvements reached when the 20% SOTA database is included in the pure synthetic database. The BPCER10 and BPCER20 for SDD-test reached 0.98% and 5.39%, respectively.